In the ideal setup shown, m1 = 3.0 kg, m2 = 2.5 kg.

a. What is the acceleration of the masses?

b. What is the tension in the string?


1
Expert's answer
2022-03-10T11:15:34-0500

a. What is the acceleration of the masses?

"a=g\frac{m_1-m_2}{m_1+m_2}"

"a=9.8*\frac{3.0-2.5}{3.0+2.5}=0.89\:\rm m/s^2"

b. What is the tension in the string?

"T=m_1g-m_1a\\
=3.0*9.8-3.0*0.89=27\:\rm N"